The pupil of a cat's eye narrows to a vertical slit of width 0.450 mm in daylight. Assume the average wavelength of the light is 615 nm. What is the angular resolution for horizontally separated mice? (In this problem, you may use the Rayleigh criterion for the limiting angle of resolution of an eye.)
rad
Answer #1

For a narrow slit, Rayleigh's criterion gives

?m = ? / a = 615 x 10-9 m /.4500 x 10-3 = 1.366 x 10-3 = 1.366 m rad
Answer #2

The same question was resolved before with different figures. Please input your figure to get your answer. Let me know if you face any issue. Please rate if I succeeded in helping you.

old question

The pupil of a cat's eye narrows to a vertical slit of width 0.450 mm in daylight. What is the angular resolution for horizontally separated mice? Assume that the average wavelength of the light is 400 nm.

Answer

Given that
width of the slit D = 0.450*10^-3m

wavelength of the light ? =500*10-9m


The angular resolution for a pair of horizontal separatedmice is
?min = 1.22 (?/D)

substiute the values.

= 1.22*500*10^-9/ 0.450*10^-3 = 1355.5555 micrometer